OpenStudy (anonymous):

The graph of f ′ (x), the derivative of x, is continuous for all x and consists of five line segments as shown below. Given f (0) = 6, find the absolute maximum value of f (x) over the interval [0, 5].

OpenStudy (anonymous):

@jim_thompson5910 Mind checking?

OpenStudy (anonymous):

The max value would have to be at that horizontal line no?

jimthompson5910 (jim_thompson5910):

They're showing f ' (x) NOT f(x)

jimthompson5910 (jim_thompson5910):

hint: find the area under the curve from x = 0 to x = 5

OpenStudy (anonymous):

So it would be 8?

OpenStudy (anonymous):

Wait, find the integral basically?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

How would I do that? Should I use geometry?

jimthompson5910 (jim_thompson5910):

yes that's how I'd do it

jimthompson5910 (jim_thompson5910):

break it up into 2 triangles and a rectangle or find the area of the trapezoid

OpenStudy (anonymous):

First triangle 2*2*1/2

OpenStudy (anonymous):

Square 2*2

OpenStudy (anonymous):

Second triangle 2*1*1/2

OpenStudy (anonymous):

Added them all.

OpenStudy (anonymous):

Which actually gives you 7

jimthompson5910 (jim_thompson5910):

so the net change is 7 meaning that f(5) = f(0) + (net change) f(5) = f(0) + 7 f(5) = 6+7 f(5) = 13 this is the max value of f(x)

OpenStudy (anonymous):

Why'd you add the 6?

jimthompson5910 (jim_thompson5910):

because we know for a fact that f(0) = 6 the graph of f ' says that f is increasing the area under f ' tells us how much f is increasing (in this case at most 7)

jimthompson5910 (jim_thompson5910):

area under f ' = net change of f(x)